# Multi-Source Urban Traffic Flow Forecasting with Drone and Loop Detector Data

### v1

@@ IV-F Partial Observation and Noisy Data
In all the experiments, HiMSNet is trained to predict the future segment speed and regional speed, which are not available from loop detectors. Therefore, the models HiMSNet(ld) and HiMSNet_PN(ld) have also benefited from the drone data in the training phase, despite that their inputs have loop detector data only. In fact, they are trained to predict future drone-measured traffic states by using past loop detector data only. This corresponds to a scenario where the training data are collected using drones, but no drones will be available after the prediction model is deployed. As a result, it is more challenging than using drone modality directly in the input, because the models need to cope with the different evolution patterns of the two data sources.
In reality, it is also possible when drone data are available neither in training nor testing, and the training labels are obtained from trajectories of taxis or ride-hailing vehicles. Although it is still possible to obtain segment speed and regional speed labels, the data quality will be much lower compared to drones, due to low penetration rates. This scenario is denoted by HiMSNet_PN(ld-), as the drone modality is considered to be absent in both training and testing phases. For implementation, the noise in training labels has zero mean and doubled standard deviation at 0.3.
TABLE II: Prediction errors with partial observation (P) and noisy (N) data
@@
Table II presents the prediction errors of HiMSNet under the partial observation and noisy data settings, best results in bold. When compared with their full-information counterparts in Table I, the PN group generally has higher prediction errors, which is expected as the model has less information and more noises to deal with.
Among all the PN settings, the model with both data modalities HiMSNet_PN has the best performance, followed by the drone-only model HiMSNet_PN(drone) and loop detector-only model HiMSNet_PN(ld). This result suggests that the drone data is a better source for urban traffic forecasting, and the loop detector data can still provide complementary information when the drone data is only partially available. However, under the full-information setting, the loop detector data shows a slightly negative effect as the single-modality setting with drone data outperforms the two-modality default setting on the segment-level task. Besides, the model with partial and noisy loop detector data HiMSNet_PN(ld) does better than its counterpart with full loop detector data HiMSNet(ld). Still, in the more realistic partial data scenario, when both modalities are present, their coverages can be complementary, and the loop detector data can provide useful information to the model and improve the prediction performance. The importance of spatial correlation manifests when the data is partial and noisy. In the segment-level task, the accuracy difference between HiMSNet_PN and HiMSNet_PN/gnn is much larger than the difference between HiMSNet and HiMSNet/gnn, which highlights the necessity of the GME module in the model. The necessity of drone modality can be further emphasized by comparing HiMSNet_PN(ld-) with HiMSNet_PN(ld), as the prediction errors have increased significantly for the regional task when the drone data is totally absent.
@@
Sensor coverage is a critical factor for prediction performance. Figure 17 shows the 30-min-ahead prediction MAE when the sensor coverage for both modalities varies simultaneously from 1% up to 100%. Firstly, it is straightforward that the higher the sensor coverage, the lower the prediction errors. Still, because of the ability to make use of correlations in space, time and modalities, HiMSNet is able to achieve similar performance for both the regional and segment-level tasks with only 20% sensor coverage. The error metrics of 10% sensor coverage make it a fair trade-off choice, if the coverage is further reduced down to 1% coverage, the performance of the model is close to the Label Average baseline in Table I. This is consistent with the intuition that when the model nearly has no information in the testing phase, the best it can do is to estimate a value based on the training samples, and such predictions (although not always constant) turn out to have nearly constant predictions and similar errors as the constant value used by Label Average.
@@ V Conclusion
In the future, this work can be improved in several directions. First, more advanced neural network architectures can be explored to better capture the complex dependencies in the traffic data, as the current model is designed to be concise and serve as a baseline. Second, the drone flight in the partial and noisy case can be carefully designed or intelligently optimized to provide best possible observation for traffic forecasting. Third, more data modalities could be included, such as vehicle flow and density. These data are also practically available and can be integrated into the model. We hope our findings can pave the way for more accurate and reliable traffic forecasting in complex urban transportation systems.

### v2

@@ IV-F Experiment Settings with Partial and Noisy Data
In all the experiments, HiMSNet is trained to predict the future segment speed and regional speed, which are not available from loop detectors. Therefore, the models HiMSNet(ld) and HiMSNet_PN(ld) have also benefited from the drone data in the training phase, despite that their inputs have loop detector data only. In fact, they are trained to predict future drone-measured traffic states by using past loop detector data only. This corresponds to a more realistic scenario where both drone and loop detector data are collected for training, but drone data are not available at the deployed environment, e.g., when drones can not fly due to adverse weather conditions. As a result, it is more challenging than using drone modality directly in the input, because the models need to cope with the different evolution patterns of the two data sources.
In reality, if drone data are not available either in training or testing, and the accurate segment speeds can only be approximately estimated using trajectories of taxis or ride-hailing vehicles. Although it is still possible to obtain segment speed and regional speed labels, the data quality will be much lower compared to drones, due to low penetration rates. This scenario is denoted by HiMSNet_PN(ld-), as the drone modality is considered to be absent in both training and testing phases. For implementation, the noise in training labels has zero mean and doubled standard deviation at 0.3.
IV-G Experiment Results with Partial and Noisy Data
TABLE III: Prediction errors with partial observation (P) and noisy (N) data
@@ IV-G Experiment Results with Partial and Noisy Data
Table III presents the prediction errors of HiMSNet under the partial observation and noisy data settings, best results in bold. When compared with their full-information counterparts in Table II, the PN group generally has higher prediction errors, which is expected as the model has less information and more noises to deal with.
The Transformer-based [33] variant HiMSNet_PN(TF) still shows very similar performance as the default HiMSNet with GCNConv [51], which again justifies the choice over a simpler design. Among all the other PN settings, the model with both data modalities HiMSNet_PN has the best performance, followed by the drone-only model HiMSNet_PN(drone) and loop detector-only model HiMSNet_PN(ld). This result suggests that the drone data is a better source for urban traffic forecasting, and the loop detector data can still provide complementary information when the drone data is only partially available. However, under the full-information setting, the loop detector data shows a slightly negative effect as the single-modality setting with drone data outperforms the two-modality default setting on the segment-level task. Besides, the model with partial and noisy loop detector data HiMSNet_PN(ld) does better than its counterpart with full loop detector data HiMSNet(ld). Although loop detector data is less effective than drone data when presented alone, its complementary coverage can still provide useful information when combined with the drone data under partial observation. The importance of spatial correlation manifests when the data is partial and noisy. In the segment-level task, the accuracy difference between HiMSNet_PN and HiMSNet_PN/gnn is much larger than the difference between HiMSNet and HiMSNet/gnn, which highlights the necessity of the GME module in the model.
While HiMSNet_PN(ld) has training labels base-on drone data, HiMSNet_PN(ld-) has no such high-quality training data. Thus HiMSNet_PN(ld-) has higher prediction errors, especially in the regional task. Nevertheless, the fact that HiMSNet_PN(ld) out performs the LA method in Table II also justifies the effects of loop detector modality. Because HiMSNet_PN(ld) is trained to predict future segment speed, and by using past point speed as hints at test time, the model can infer the segment speed better than taking an average over historical data.
@@
Sensor coverage is a critical factor for prediction performance. Figure 17 shows the 30-min-ahead prediction MAE when the sensor coverage for both modalities varies simultaneously from 1% up to 100%. Firstly, it is straightforward that the higher the sensor coverage, the lower the prediction errors. Still, because of the ability to make use of correlations in space, time and modalities, HiMSNet is able to achieve similar performance for both the regional and segment-level tasks with only 20% sensor coverage, when comapred to the full coverage case. The error metrics of 10% sensor coverage make it a fair trade-off choice, if the coverage is further reduced down to 1% coverage, the performance of the model is close to the Label Average baseline in Table II. This is consistent with the intuition that when the model nearly has no information in the testing phase, the best it can do is to estimate a value based on the training samples, and such predictions (although not always constant) turn out to have nearly constant predictions and similar errors as the constant value used by Label Average.
@@ V Conclusion
In the future, this work can be improved in several directions. First, more advanced neural network architectures can be explored to better capture the complex dependencies in the traffic data, as the current model is designed to be concise and serve as a baseline. Second, the drone flight in the partial and noisy case can be carefully designed or intelligently optimized to provide best possible observation for traffic forecasting. Third, more data modalities could be included, such as vehicle flow and density. These data are also practically available and can be integrated into the model. Besides, real-world experiments can be conducted to integrate the prediction approach with real-time traffic data platforms and develop deliverable services for traffic management. Furthermore, hierarchical clustering can reduce the computational cost of the model and facilitate the deployment of the model in real-world applications. Finally, we hope our findings can pave the way for more accurate and reliable traffic forecasting in complex urban transportation systems.